# N2: A Unified Python Package and Test Bench for Nearest Neighbor-Based Matrix Completion

### v1

@@ D.3 MovieLens
We load MovieLens via a custom MovieLensDataLoader that (i) downloads and caches the ml-1m.zip archive, (ii) reads ratings.dat into a user $\times$ movie pivot table, and (iii) constructs the binary mask where observed entries correspond to rated user–movie pairs. The data matrix is $Z\in\{1,\dots,5\}^{6040\times 3952}$ and mask matrix is $A\in\{0,1\}^{6040\times 3952}$. The data can be downloaded from https://grouplens.org/datasets/movielens/1m/. See https://files.grouplens.org/datasets/movielens/ml-1m-README.txt for the usage license.
@@ D.4 Proposition 99
Data comes primarily from the Tax Burden on Tobacco compiled by Orzechowski and Walker [41] (ODC-By License). Using synthetic control methods, Abadie et al. construct a weighted combination of control states that closely resembles California’s pre-1988 characteristics and cigarette consumption patterns. The optimal weights produce a synthetic California primarily composed of Colorado (0.164), Connecticut (0.069), Montana (0.199), Nevada (0.234), and Utah (0.334), with all other states receiving zero weight. The treatment effect is estimated as the difference between actual California per-capita cigarette sales and those of synthetic California after Proposition 99’s implementation. By 2000, this analysis revealed that annual per-capita cigarette sales in California were approximately 26 packs lower than what they would have been without Proposition 99, representing about a 25% reduction in cigarette consumption. To validate these findings, the authors conducted placebo tests by applying the same methodology to states not implementing tobacco control programs, confirming that California’s reduction was unusually large and statistically significant (p = 0.026).
Proposition 99, the California Tobacco Tax and Health Protection Act of 1988, dataset spans from 1970 to 2000, providing 19 years of pre-intervention data before Proposition 99 was implemented in 1988 and 12 years of post-intervention data. It provides annual state-level cigarette consumption measured as per capita cigarette sales in packs based on tax revenue data. This data serves as a real data benchmark for many of the variants of synthetic controls [2]. We use the CDC dataset for the Nearest Neighbors methods and only use the target variable (i.e., cigarette consumption measured in packs per capita), and the dataset from SyntheticControlMethods library [^3] https://github.com/OscarEngelbrektson/SyntheticControlMethods/tree/master (Apache-2.0 License) for the SC baseline, since it relies on additional covariates.
@@ D.5 PromptEval
MMLU is a multiple choice Q&A benchmark with $57$ tasks, with a total of near $14$K examples [^4] https://github.com/felipemaiapolo/prompteval (MIT License). $15$ different models, e.g., variants of Llama 3 [39], Mistral [31] and Gemma [50]. The examples are fed into the models with $100$ different varying prompting templates. The prompt templates are created by traversing between $3$ node modules, namely a separator, a space and an operator (see [43, Alg. 3, App. J] for details), from which $100$ unique prompt templates are created. The unitxt [12] preprocessing library is used to construct the dataset and evaluation is done by LM-Eval-Harness [26] library. The number of examples differ per task and each examples are evaluated on a model (verifiable, so assigned 00 or $1$ for correctness) by wrapping the examples with $100$ different prompt templates.

### v2

@@ D.3 MovieLens
We load MovieLens via a custom MovieLensDataLoader that (i) downloads and caches the ml‑1m.zip archive, (ii) reads ratings.dat into a user $\times$ movie pivot table, and (iii) constructs the binary mask where observed entries correspond to rated user–movie pairs. The data matrix is $Z\in\{1,\dots,5\}^{6040\times 3952}$ and mask matrix is $A\in\{0,1\}^{6040\times 3952}$. The data can be downloaded from https://grouplens.org/datasets/movielens/1m/. See https://files.grouplens.org/datasets/movielens/ml-1m-README.txt for the usage license.
@@ D.4 Proposition 99
Data comes primarily from the Tax Burden on Tobacco compiled by Orzechowski and Walker [cdcBurdenTobacco] (ODC-By License). Using synthetic control methods, Abadie et al. construct a weighted combination of control states that closely resembles California’s pre-1988 characteristics and cigarette consumption patterns. The optimal weights produce a synthetic California primarily composed of Colorado (0.164), Connecticut (0.069), Montana (0.199), Nevada (0.234), and Utah (0.334), with all other states receiving zero weight. The treatment effect is estimated as the difference between actual California per-capita cigarette sales and those of synthetic California after Proposition 99’s implementation. By 2000, this analysis revealed that annual per-capita cigarette sales in California were approximately 26 packs lower than what they would have been without Proposition 99, representing about a 25% reduction in cigarette consumption. To validate these findings, the authors conducted placebo tests by applying the same methodology to states not implementing tobacco control programs, confirming that California’s reduction was unusually large and statistically significant (p = 0.026).
Proposition 99, the California Tobacco Tax and Health Protection Act of 1988, dataset spans from 1970 to 2000, providing 19 years of pre-intervention data before Proposition 99 was implemented in 1988 and 12 years of post-intervention data. It provides annual state-level cigarette consumption measured as per capita cigarette sales in packs based on tax revenue data. This data serves as a real data benchmark for many of the variants of synthetic controls [athey2021matrix]. We use the CDC dataset for the Nearest Neighbors methods and only use the target variable (i.e., cigarette consumption measured in packs per capita), and the dataset from SyntheticControlMethods library [^4] https://github.com/OscarEngelbrektson/SyntheticControlMethods/tree/master (Apache-2.0 License) for the SC baseline, since it relies on additional covariates.
@@ D.5 PromptEval
MMLU is a multiple choice Q&A benchmark with $57$ tasks, with a total of near 1414K examples [^5] https://github.com/felipemaiapolo/prompteval (MIT License). $15$ different models, e.g., variants of Llama 3 [llama3], Mistral [jiang2023mistral] and Gemma [team2024gemma]. The examples are fed into the models with $100$ different varying prompting templates. The prompt templates are created by traversing between $3$ node modules, namely a separator, a space and an operator (see [polo2024efficient, Alg. 3, App. J] for details), from which $100$ unique prompt templates are created. The unitxt [bandel2024unitxt] preprocessing library is used to construct the dataset and evaluation is done by LM-Eval-Harness [eval-harness] library. The number of examples differ per task and each examples are evaluated on a model (verifiable, so assigned 0 or $1$ for correctness) by wrapping the examples with $100$ different prompt templates.